# Quantifying $S_{8}$ tension and evidence for interacting dark energy from redshift-space distortion measurements

### v1

@@ IV.1 Constraints without CMB data
In Table 1, we summarize the results for the case where $\xi<0$. As anticipated, RSD data alone do not possess the capability to robustly restrict the entire parameter space of the model. However, several noteworthy points emerge. As expected, the RSD data are not sensitive to the Hubble constant, which prevents us from obtaining precise constraints on $H_{0}$ from RSD data alone. The constraints on $\Omega_{\rm m}$ and $S_{8}$ show significantly lower values compared to those predicted by the $\Lambda$CDM model, albeit with large error bars. However, the most interesting observation is that the coupling parameter $\xi$ is well-constrained using RSD data alone. Specifically, we find $\xi>-0.0331$ at 95% CL, indicating that RSD can impose strong bounds on the total amount of energy-momentum that can be transferred from DM to DE. Notably, in this case, RSD data improve the constraints on the coupling parameter by an order of magnitude compared to CMB data alone (see CMB-only constraints in Zhai et al. (2023b)).
@@
As a next step, we perform a joint analysis of RSD+DESI+PP data. We obtain tight constraints on the entire parameter space of the model. Specifically, the coupling parameter is significantly constrained to $\xi>-0.0133$ at 95% CL. The parameter $S_{8}$ is robustly constrained to $S_{8}=0.773\pm 0.025$ at 68% CL, indicating potentially low values of $S_{8}$ from the combination of RSD+DESI+PP data. However, we do not observe any evidence supporting a non-zero coupling parameter, as it is tightly constrained by the joint analysis. As quantified in Table 1 and Figure 2, the addition of CC data does not result in significant statistical improvements. Therefore, RSD+DESI+PP alone is robust enough to constrain the entire parameter space of this model.
Now, we turn our attention to the scenario where $\xi>0$. We summarize the results for this case in Table 2. When considering RSD data, we observe behavior opposite to the $\xi<0$ case. This reflects the opposite direction in which the energy-momentum is transferred within the dark sector (here from DE to DM), introducing opposite correlations between the coupling $\xi$ and the other baseline parameters. Specifically, the constraints on the parameters $\Omega_{\rm m}$, $H_{0}$, and $S_{8}$ move in the opposite direction. With $\xi>0$, RSD data tend to predict higher values for both $\Omega_{\rm m}$ and $S_{8}$, although the error bars remain large since RSD data alone lack strong constraining power. Conversely, $H_{0}$ shifts towards lower values with error bars improving by up to a factor of 2. Overall, when $\xi>0$, RSD data on their own do not provide robust constraints on the free parameters of the model, including $\xi$.
@@
Next, we pursue the analysis with the combined data from RSD+DESI+PP. In this case, we find $\xi>0$ with more than $2\sigma$ confidence. This joint analysis robustly constrains the full parameter space of the model. We obtain $H_{0}=68.99\pm 0.68$ km/s/Mpc, which is in tension with local measurements by the SH0ES team. In contrast to the $\xi<0$ case, we observe a slightly higher value for $S_{8}$. This is due to the strong correlation between $\xi$ and $\sigma_{8}$, which acts in opposite directions depending on the sign of $\xi$.
Figure 3 shows the one-dimensional and two-dimensional marginalized posterior distributions (68% and 95% CL) for the parameters of interest in the model, based on our statistical analyses. It is evident from these contours that the inclusion of RSD data significantly enhances the constraining power on the parameters, leading to a more precise determination of their values and reducing uncertainties. Notably, the linear perturbation effects captured by RSD measurements show a preference for $\xi>0$, which also explains the strong limit on $\xi<0$ since the data favor a positive coupling. Additionally, the inclusion of CC data does not affect the results for $\xi>0$, reinforcing that RSD+PP+DESI alone provides robust constraints on the model parameters.
@@ IV.2 Constraints with CMB data
To avoid combining datasets with significant tension that could possibly bias results for crucial parameters, such as the coupling parameter quantifying energy-momentum transfer within the dark sector (which, as our CMB-free analysis shows, is highly correlated with parameters exhibiting the greatest tension), we will focus on the case $\xi>0$. In Table 3, we summarize our statistical results, showing the constraints at 68% and 95% CL for our baseline parameters of interest. When analyzing CMB data only, we find an upper limit for the coupling parameter $\xi<0.388$ at 95% CL. As expected, the strong correlation between $\xi$ and the other parameters (see Figure 5) drives $H_{0}$ and $S_{8}$ towards lower values compared to the case where $\xi<0$. As a result, while IDE models with energy-momentum transfer from DE to DM do not alleviate the Hubble tension, they can significantly help address (or at least relax) the tension in $S_{8}$. The predicted $S_{8}$ value is low enough to be fully compatible with the RSD-only analysis discussed in the previous section, as well as with cosmic shear measurements. Again, this is clearly illustrated in the right panel of Figure 4, where the probability contours for $S_{8}$ and $\Omega_{m}$ inferred from the two probes overlap well within the 68% CL region of the parameter space. This comprehensive analysis underscores the importance of considering the correlations between parameters and highlights how certain scenarios can effectively reconcile discrepancies between different observational datasets.
@@
As a next step, we consider different combinations with DESI and PP data. For the CMB+RSD+DESI combination, the evidence for $\xi$ is found exactly at $2\sigma$. For the CMB+RSD+DESI+PP+CC combination, the evidence for $\xi>0$ is again greater than $2\sigma$ CL. In our most robust joint analysis, i.e., CMB+RSD+DESI+PP+CC, we find $\xi=0.029^{+0.025}_{-0.026}$ at 95% CL, and $S_{8}=0.7950\pm 0.0091$ at 68% CL. We refer to Figure 5 for the one- and two-dimensional marginalized probability distributions of the parameters $\Omega_{\rm m}$, $\xi$, $H_{0}$, and $S_{8}$ obtained from all these data combinations.
@@ V Final Remarks
  • New Bounds for Dark Coupling: RSD data has a strong impact on models with $\xi<0$, allowing us to impose very significant upper bounds in these cases. Specifically, we find that $\xi>-0.0331$ from RSD data alone. Combination with PP and DESI samples, further improves this bounds to $\xi>-0.0133$.
  • Solution for the $S_{8}$ Tension: Due to the strong correlation between $\xi$ and other critical parameters such as $\Omega_{m}$, $H_{0}$, and $S_{8}$ (see also Figure 3), positive values of $\xi$ lead to lower values of $H_{0}$ and $S_{8}$ compared to the case where $\xi<0$. Therefore, while IDE models featuring energy-momentum transfer from DE to DM do not alleviate the Hubble tension, they effectively resolve the tension between CMB and RSD data regarding $\Omega_{m}$ and $S_{8}$, as shown in the right panel of Figure 4. This highlights the plausibility of models with $\xi>0$ in achieving values of $S_{8}$ that align with both datasets, thereby addressing the $S_{8}$ tension.
[…]
@@ Acknowledgements.
M.A.S and E.S. received support from the CAPES scholarship. R.C.N. thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnologico (CNPq, National Council for Scientific and Technological Development) under the project No. 304306/2022-3, and the Fundação de Amparo à Pesquisa do Estado do RS (FAPERGS, Research Support Foundation of the State of RS) for partial financial support under the project No. 23/2551-0000848-3. S.K. gratefully acknowledges the support of Startup Research Grant from Plaksha University (File No. OOR/PU-SRG/2023-24/08), and Core Research Grant from Science and Engineering Research Board (SERB), Govt. of India (File No. CRG/2021/004658). E.D.V. is supported by a Royal Society Dorothy Hodgkin Research Fellowship. W.G. is supported by the Lancaster–Sheffield Consortium for Fundamental Physics under STFC grant: ST/X000621/1. This article is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology).

### v2

@@ IV.1 Constraints without CMB data
In Table 1, we summarize the results for the case where $\xi<0$. As anticipated, RSD data alone do not possess the capability to robustly restrict the entire parameter space of the model. However, several noteworthy points emerge. As expected, the RSD data are not sensitive to the Hubble constant, which prevents us from obtaining precise constraints on $H_{0}$ from RSD data alone. The constraints on $\Omega_{\rm m}$ and $S_{8}$ show significantly lower values compared to those predicted by the $\Lambda$CDM model, albeit with large error bars. However, the most interesting observation is that the coupling parameter $\xi$ is well-constrained using RSD data alone. Specifically, we find $\xi>-0.0440$ at 95% CL, indicating that RSD can impose strong bounds on the total amount of energy-momentum that can be transferred from DM to DE. Notably, in this case, RSD data improve the constraints on the coupling parameter by an order of magnitude compared to CMB data alone (see CMB-only constraints in Zhai et al. (2023b)).
@@
As a next step, we perform a joint analysis of RSD+DESI+PP data. We obtain tight constraints on the entire parameter space of the model. Specifically, the coupling parameter is constrained to $\xi>-0.0430$ at 95% CL. The parameter $S_{8}$ is robustly constrained to $S_{8}=0.775\pm 0.027$ at 68% CL, indicating potentially low values of $S_{8}$ from the combination of RSD+DESI+PP data. However, we do not observe any evidence supporting a non-zero coupling parameter, as it is tightly constrained by the joint analysis. As quantified in Table 1 and Figure 2, the addition of CC data does not result in significant statistical improvements. Therefore, RSD+DESI+PP alone is robust enough to constrain the entire parameter space of this model. These analyses clearly demonstrate that RSD data improve the constraints significantly on this class of models.
Now, we turn our attention to the scenario where $\xi>0$. We summarize the results for this case in Table 2. When considering RSD data, we observe behavior opposite to the $\xi<0$ case. This reflects the opposite direction in which the energy-momentum is transferred within the dark sector (here from DE to DM), introducing opposite correlations between the coupling $\xi$ and the other baseline parameters. Specifically, the constraints on the parameters $\Omega_{\rm m}$, $H_{0}$, and $S_{8}$ move in the opposite direction. With $\xi>0$, RSD data tend to predict higher values for both $\Omega_{\rm m}$ and $S_{8}$, although the error bars remain large since RSD data alone lack strong constraining power. Overall, when $\xi>0$, RSD data on their own do not provide robust constraints on the free parameters of the model, including $\xi$.
@@
Next, we pursue the analysis with the combined data from RSD+DESI+PP. In this case, we find $\xi<0.0717$ at 95% CL. This joint analysis robustly constrains the full parameter space of the model. We obtain $H_{0}=69.37\pm 0.78$ km/s/Mpc, which is in tension with local measurements by the SH0ES team. In contrast to the $\xi<0$ case, we observe a slightly higher value for $S_{8}$. This is due to the strong correlation between $\xi$ and $\sigma_{8}$, which acts in opposite directions depending on the sign of $\xi$.
Figure 3 shows the one-dimensional and two-dimensional marginalized posterior distributions (68% and 95% CL) for the parameters of interest in the model, based on our statistical analyses.
It is evident from these contours that the inclusion of RSD data significantly enhances the constraining power on the parameters, leading to a more precise determination of their values and a reduction in uncertainties. Notably, the linear perturbation effects captured by RSD measurements impose stronger upper bounds on the coupling parameter $\xi$ in this context. Additionally, the inclusion of CC data does not impact the results for $\xi>0$, further reinforcing that RSD+PP+DESI alone provides robust constraints on the model parameters.
@@ IV.2 Constraints with CMB data
To avoid combining datasets with significant tension that could possibly bias results for crucial parameters, such as the coupling parameter quantifying energy-momentum transfer within the dark sector (which, as our CMB-free analysis shows, is highly correlated with parameters exhibiting the greatest tension), we will focus on the case $\xi>0$. In Table 3, we summarize our statistical results, showing the constraints at 68% and 95% CL for our baseline parameters of interest. When analyzing CMB data only, we find an upper limit for the coupling parameter $\xi<0.388$ at 95% CL. As expected, the strong correlation between $\xi$ and the other parameters (see Figure 5) drives $H_{0}$ and $S_{8}$ towards lower values compared to the case where $\xi<0$. As a result, while IDE models with energy-momentum transfer from DE to DM do not alleviate the Hubble tension, they can significantly help address (or at least relax) the tension in $S_{8}$. The predicted $S_{8}$ value is low enough to be fully compatible with the RSD-only analysis discussed in the previous section, as well as with cosmic shear measurements. Again, this is clearly illustrated in the right panel of Figure 4, where the probability contours for $S_{8}$ and $\Omega_{m}$ inferred from the two probes overlap well within the 95% CL region of the parameter space. This comprehensive analysis underscores the importance of considering the correlations between parameters and highlights how certain scenarios can effectively reconcile discrepancies between different observational datasets.
In the joint CMB+RSD analysis, we observe evidence for an energy transfer from DE to DM at 68% CL, although this effect is consistent with zero at 95% CL. Specifically, we find that $\xi<0.0545$ at 95% CL in this joint analysis. This result is noteworthy, clearly indicating that the combination of evolution-dependent observables from linear perturbations of matter with CMB leads to an improvement in the upper limits when compared to BAO and SNe Ia data.
As a next step, we explore different combinations involving DESI and PP data. The inclusion of these observations not only improves the precision of the constraints but also results in slightly lower $S_{8}$ values. For the CMB+RSD+DESI combination, we find an upper limit of $\xi<0.0405$. In the case of the CMB+RSD+DESI+PP+CC combination, the upper limit shifts to $\xi<0.0449$. The slight increase in the upper limit of $\xi$ is driven by the influence of SNIa data, which tends to pull $\xi$ towards more positive values. It is evident that RSD predictions play a crucial role in constraining the parameter space for this class of models, resulting in such tight upper limits. We refer to Figure 5 for the one- and two-dimensional marginalized probability distributions of the parameters $\Omega_{\rm m}$, $\xi$, $H_{0}$, and $S_{8}$ obtained from all these data combinations.
@@
At this stage, it is useful to compare our results for $S_{8}$ with the current weak lensing and galaxy clustering measurements predicted by surveys such as DES Abbott et al. (2022) and KiDS-1000 Asgari et al. (2021). From our analysis based on CMB data alone, we find $S_{8}=0.787\pm 0.024$, a value that is slightly low and consistent with the cosmic shear measurements from DES and KiDS-1000. When considering our joint analysis, which incorporates CMB, RSD, and DESI data, we find $S_{8}=0.796\pm 0.010$, which exhibits a tension of approximately $1.5\sigma$ with KiDS-1000, and a similar degree of tension with the DES results. Since these models modify the late-time growth of cosmic structures, a more quantitative comparison requires re-analyzing the weak lensing and galaxy clustering data from these surveys, as the constraints from the cosmic shear samples depend on the specific cosmological model. In future work, we will explore this issue in more detail to address the model-dependence of these comparisons.
@@
In all the tests we conducted, we observed that measurements at low redshift values (for $z<0.5$) possess a significant capability to differentiate between IDE and the standard $\Lambda$CDM model. This finding highlights the critical importance of low-redshift data in constraining the parameters of these cosmological models. Consequently, other RSD compilations that incorporate additional data within this redshift range may yield different results than those discussed above. Such variations could lead to new insights into the nature of DE and its interactions, emphasizing the need for a comprehensive analysis of low-redshift observations in future studies.
@@ V Final Remarks
  • New Bounds for Dark Coupling: RSD data has a strong impact on models with $\xi<0$, allowing us to impose very significant upper bounds in these cases. Specifically, we find that $\xi>-0.0440$ from RSD data alone. Combination with PP and DESI samples, further improves this bounds to $\xi>-0.0430$. Similar conclusions are drawn for $\xi>0$, where we obtain the upper limit $\xi<0.0716$ from the combination of DESI, PP, RSD, and CC data.
  • Solution for the $S_{8}$ Tension: Due to the strong correlation between $\xi$ and other key cosmological parameters such as $\Omega_{m}$, $H_{0}$, and $S_{8}$ (see also Figure 3), positive values of $\xi$ suppress the growth factor at low redshift ($z<0.75$), while maintaining $\Omega_{m}\sim 0.31$, resulting in lower $S_{8}$ values compared to the $\Lambda$CDM scenario. Although IDE models featuring energy-momentum transfer from DE to DM do not resolve the Hubble tension, they do effectively mitigate the tension between CMB and RSD data concerning $\Omega_{m}$ and $S_{8}$, as shown in the right panel of Figure 4. This underscores the viability of models with $\xi>0$ in achieving $S_{8}$ values that reconcile both datasets, thereby addressing the $S_{8}$ tension.
  • Interaction Evidence: Based on the statistical criteria used to select the RSD sample, we found that RSD data have a significant impact on this class of models, imposing stringent upper limits on the parameter $\xi$ across all cases analyzed in this work. Consequently, we conclude that the inclusion of RSD data strongly constrains IDE models, rendering them nearly indistinguishable from the standard $\Lambda$CDM model.
@@ Acknowledgements.
The authors express their gratitude to the referee for the valuable comments and suggestions, which have contributed to enhancing the significance of the results presented in this work. M.A.S and E.S. received support from the CAPES scholarship. R.C.N. thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnologico (CNPq, National Council for Scientific and Technological Development) under the project No. 304306/2022-3, and the Fundação de Amparo à Pesquisa do Estado do RS (FAPERGS, Research Support Foundation of the State of RS) for partial financial support under the project No. 23/2551-0000848-3. S.K. gratefully acknowledges the support of Startup Research Grant from Plaksha University (File No. OOR/PU-SRG/2023-24/08), and Core Research Grant from Science and Engineering Research Board (SERB), Govt. of India (File No. CRG/2021/004658). E.D.V. is supported by a Royal Society Dorothy Hodgkin Research Fellowship. W.G. is supported by the Lancaster–Sheffield Consortium for Fundamental Physics under STFC grant: ST/X000621/1. This article is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology).